Circular No. 8659

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN J. A. Farrell, Jemez Springs, NM, reports that unfiltered CCD expo-

sures obtained with a 0.41-m f/8.4 reflector show a companion to comet 73P moving at the same general direction and rate as the primary; separations and position angles (± 0^{o}.02) from the primary are as follows: Jan. 6.48831 UT, 1392′′.3 ± 0′′.6 in p.a. 270^{o}.28; 7.51252, 1414′′.4 ± 0′′.8 in p.a. 269^{o}.90; 9.39304 (through high clouds), 1454′′.5 ± 1′′.8 in p.a. 270^{o}.04. On the first two nights, the magnitude of the primary was found to be 16.2–16.4, while the secondary was at mag 18.8–19.0.

Z. Sekanina, Jet Propulsion Laboratory, provides a search ephemeris for the companion discovered by Farrell. The few data points available can be reasonably well linked with the 1995–1996 observations of nucleus B, but the final identification should wait for additional observations in coming weeks. The newly found companion is not identical with fragments A, E, or F observed between 1995 and 2001 (cf. Sekanina 2005, ICQ 27, 225-

240; online at http://cfa-www.harvard.edu/icq/sek icq V27 225.pdf).

The predicted separation distances from the main comet (in arcmin) and position angles (in degrees, for equinox J2000.0) are as follows: 2006 Jan. 15.0 TT, 29′.9, 297^{o}.9; 25.0, 35′.4, 298^{o}.0; Feb. 4.0, 42′.3, 297^{o}.9; 14.0, 51′.3, 297^{o}.6; 24.0, 63′.0, 297^{o}.2; Mar. 6.0, 78′.6, 296^{o}.5.

COMET P/2005 YQ_{127} (LINEAR) Further to IAUC 8658, Mannucci et al. note the presence of a faint 7′′

coma on their images from Jan. 9.9 UT, and numerous other CCD observations confirming the cometary nature of this object have been reported. F. Bellini and L. Buzzi (Varese, Italy, 0.60-m reflector) write that the object was clearly diffuse with a moderate central condensation and a round coma ∼ 10′′ in diameter, elongated in p.a. 135^{o} on Jan. 10.05. S. Foglia, D. Crespi, G. Galli, and S. Minuto (near Novara Veveri, Italy, 0.40-m f/4 reflector) note that images taken on Jan. 9.91 show a slightly diffuse coma of diameter 7′′ elongated in p.a. 240^{o}. P. Corelli (Pagnacco, Italy, 0.20-m Schmidt-Cassegrain reflector) adds that images taken on Jan. 8.76 and 9.80 show a faint, diffuse coma of diameter 15′′. C. W. Hergenrother (Steward Observatory, 1.54-m Kuiper telescope) found a diffuse 40′′ coma on his 120sec R-band exposures taken on Jan. 4.39. Images taken by E. Guido and G. Sostero (Castellammare di Stabia, Italy, 0.25-m f/3.4 reflector) on Jan. 10.3 show a faint, diffuse coma nearly 20′′ with a total R-band magnitude of ∼ 16.0.

2006 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green